Solve the system of equations by graphing.
x+y=4
y-x=4
Please show points, how to graph the points and the point at where they meet or intersect (if they do).

Answer by Alan3354(69443)

Solve the system of equations by graphing.
x+y=4
y-x=4
----------
It's 2 straight lines.
x+y=4
Find 2 points. Pick values for x or y and find the other variable. Zero usually works well.
Plot the 2 points.
Draw a line thru them.
-------------
Do the same for the other equation.
Where they intersect is at the values of x & y that fit both equations.